OpenStudy (anonymous):

a particle starts SHM from means postition then find out the time taken by particle in moving from mean position to half of extreme position and from half of extreme to mean position

OpenStudy (wasiqss):

so

OpenStudy (anonymous):

my doubt is how to start i mean in first part x=asinwt

OpenStudy (anonymous):

no shankvee its not the ans by the way and u dont jhave to solve till i ask u ..ok u just have to make correction in my ans or clear my doubts dat i m facing in ques

OpenStudy (anonymous):

ok ok very sorry. :D

OpenStudy (anonymous):

yeah so in first part half of extreme would be A/2. Then find time.

OpenStudy (anonymous):

np.. well here extreme means A rit and half of extreme is said to be A/2 so A/2=Asin wt sinwt=/2 sinwt=pie/6

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

nw from extreme to mean i m not gettin how to do this part

OpenStudy (anonymous):

i mean give me hint dont solve

OpenStudy (anonymous):

Whatever time it takes from mean to extreme its same for reverse journey.

OpenStudy (anonymous):

no its not and how can be it will be same imagine there is air ressitance force is also there which slows down the speed

OpenStudy (anonymous):

i remeber when i used pendulum timing is not same in practicals

OpenStudy (anonymous):

Ok lets see (Air resistance does not matter This is ideal case.)

OpenStudy (anonymous):

as velocity decreases due to time increases

OpenStudy (wasiqss):

shm means that we have to neglect resistive force

OpenStudy (anonymous):

Ok so if you calculate time to reach from mean to extreme then you get wt=pi/2. Now time taken to reach from mean to mean again, would be wt=pi. Think about it. subtracting the two would give wt=pi/2 so it turns out to be same.

OpenStudy (anonymous):

but then dont u think pendulum will have same KE in all position??

OpenStudy (anonymous):

No KE isn't same as PE varies( Conservation of mechanical energy). Total mechanical energy will reamin same.

OpenStudy (anonymous):

thats wat i m saying velocity differs due t which KE also differs

OpenStudy (anonymous):

to

OpenStudy (anonymous):

Wait what are you asking?

OpenStudy (anonymous):

time period from half of extreme to mean

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Well here  Now here particle will first move from half of mean to extreme and then from extreme to mean. Calculate each time seperately.